OpenStudy (anonymous):

PLEASE HELP!!!!! The sum of two consecutive even integers is 234. What is the larger integer? 114 116 118 124

OpenStudy (perl):

x + (x+2) = 234 2x + 2 = 234 2x = 232 x = 116 . so the larger integer is 118

OpenStudy (zehanz):

You will have to find a way to describe these numbers. To do so, first think how you can describe an even number: You could call it 2a, because it is even, so you can divide it by 2. This means 2 is a factor, so the number is 2*another number, or 2a. Now the next even number is 2 more than 2a, so it is 2a+2. We can change the whole story into an equation now: 2a(2a+2)=118. If you solve it, you have got a. Then the larger number is 2a+2. Will you try this now?

OpenStudy (anonymous):

Thank you @perl :)

OpenStudy (anonymous):

Ok. So the answer is 118?

OpenStudy (zehanz):

@perl: it would be better to leave something to do for RosieF, instead of just giving the complete answer IMO.

OpenStudy (zehanz):

Then there is the typo in my reply: 2a + 2a+2 = 234. This now has to be solved.

OpenStudy (anonymous):

Thank you 2 @ZeHanz :) I'd give u a medal 2 if I could ;)

OpenStudy (perl):

@ZeHanz it is not necessary to write it the integer as 2a

OpenStudy (perl):

my method is straightforward, it works for even and odd integer consecutive problems

OpenStudy (zehanz):

@perl I know, the numbers being even is irrelevant. But only afterwards that becomes clear.